OpenStudy (anonymous):

What is the equation of the line that passes through the points (-1, 3) and (1, 11)? 4x - y = 1 4x + y = 1 x - 4y = -13 4x - y = -7

OpenStudy (abdullah_me):

4x - y = -7 because slope =4 equation y-3=4(x-(-1))

OpenStudy (anonymous):

thank you so much :)